Stinler Corporation manufactures curtains for retail stores and for individual customers. The retail chain buys 70% of the curtains produced. The individual customers purchase curtains in approximately equal quantities, where the orders are about the same size. Data concerning Stinler's customer activity are as follows:
Retail Stores Individual Customers
Units sold 56,000 24,000
Orders placed 8 492
Number of sales calls 4 156
Manufacturing costs 8,000,000 3,500,000
Order filling costs allocated 375,000 225,000
Sales force costs allocated 265,000 135,000
Currently, allocations of Order Filling and Sales Force costs are assigned to customers based on units sold, a unit-level driver.
Show all calculation steps to
Re-assign Order Filling and Sales Force costs to customers by using an ABC approach.
Answer #1

Under ABC costing, overhead costs are divided into different costs and are assigned using different cost drivers

Order filling costs will be allocated as per number of orders placed and sales force cost shall be allocated using number of sales calls

Retail Stores

Individual Customers

Order filling costs allocated

600,000*8/500 = $9,600

600,000*492/500 = $590,400

Sales force costs allocated

400,000*4/160 = $10,000

400,000*156/160 = $390,000

Total cost allocated

19,600

980,400